OpenStudy (anonymous):

how do I sketch y = 1 + ln(x-2)?

OpenStudy (anonymous):

Just like any other graph. Find values for x

OpenStudy (anonymous):

how about in term of shifts: y = x^2 + 5, shift y = x^2 up by 5

OpenStudy (anonymous):

that was just an example. I was asking what I need to do to the graph of y = ln x to get the graph I want.

OpenStudy (anonymous):

ln x is a number in the same way x^2 is a number.

OpenStudy (anonymous):

Not quite, the domain is restricted to positive reals.

OpenStudy (anonymous):

I am not saying they are like or similar numbers; but he or she is unsure about ln x: I am merely saying it is a number.

OpenStudy (anonymous):

I am saying that for sketching purposes x must be >2.

OpenStudy (anonymous):

ok, x must be greater than 2, what about the 1, do I shift the graph upward by 1?

OpenStudy (anonymous):

just choose any numbers applicable and get an idea of the graph and draw it. There is no need to shift, unless that is some specific instructions your teacher gave you.

OpenStudy (anonymous):

Choose values of x higher than 2.

OpenStudy (anonymous):

If you set x = 3, y = 1 + ln 1 = 1

OpenStudy (anonymous):

The normal graph of ln is asymptotic to negative y....